OpenStudy (anonymous):

An object moving along a curve in the xy-plane has position (x(t),y(t)) at time t>=0 with dx/dt=3+cos(t^2). The derivative dy/dt is not explicitly given. At time t=2, the object is at position (1,8). (a) Find the x-coordinate of the position of object at time t=4. (b) At time t=2, the value of dy/dt is -7. Write an equation for the line tangent to the curve at the point (x(2),y(2)) (c) Find the speed of the object at time t=2. (d) For t>=3, the line tangent to the curve at (x(t),y(t)) has a slope of 2t+1. Find acceleration vector of the object at time t=4.

OpenStudy (anonymous):

Please give me hints rather than the entire solution.

OpenStudy (experimentx):

a) integrate 3+cos(t^2) dt ... use initial value of x(2) = 1

OpenStudy (anonymous):

I have to integrate it?

OpenStudy (experimentx):

yes .. integrate if form 2 to 4

OpenStudy (anonymous):

I don't know how to integrate cos(t^2)

OpenStudy (experimentx):

are you allowed to use calculator?? i mean some sorta graphic calculator or .. wolfram?? this is definite integral ... my guess is ... indefinite integral is not closed or non elementary http://www.wolframalpha.com/input/?i=integrate+cos%28x^2%29

OpenStudy (experimentx):

add 1 to this http://www.wolframalpha.com/input/?i=integrate+cos%28x^2%29%2B3+from+2+to+4

OpenStudy (experimentx):

b) ... calculate the dx/dt at t=2 .. ie. dx/dt=3+cos(t^2)|t=2 dy/dx = dy/dt* 1/(dx/dt) <=== this slope m, since you know coordinate 1,8 find tangent using single point formula

OpenStudy (experimentx):

dy/dt is given, dx/dt ... you calculate in step b) ds/dt = \( \sqrt{(dx/dt)^2 + (dy/dt)^2}\)

OpenStudy (experimentx):

d) the acceleration is given by d^2s/dt^2 dy/dx = 2t+1 dy/dt = (2t+1) dx/dt use this formula to calculate ds/dt differentate it again ...wrt t and put t=4 this will give your acceleration

OpenStudy (anonymous):

Okay now I know why he only gave two questions for homework.... It is super difficult!! Thank you, @experimentX !

OpenStudy (experimentx):

well ... :) good night!!